and trees from invading and becomine establishedThe be thick and nutrient rich, nnd the roots of perennint plants are very deep. Blome present in n Olcinhoms, central Ain, etc.. 22. This biome is characterized by wet, mild wintern and dry summers Fire-resistant, evergreen, woody nhrubs dominnte. Biome present Califormin, South Afien, etc 29. Which of the following is a density-dependent factor limiting the growth of populations? A. Space on rocky outcrops for mussels to settle B. Damage caused by floating logs to mussel beds C. Compctition for light among redwood seedlings D. A and C E. All of the above Question 30 refers to a community that contains the following species in the following numbers: Species Species A Species C Number 16 Species B Species D 24 30. What is the diversity index of this Community? A. 5 B. .613 C. .387 D. 100 E. .0325 31. Character displacement A. Is the result of intraspecific competition B. Is the result of interspecific competition C. Causes sympatric populations of certain species to differ in body structure from allopatric populations D. Results in the extinction of the inferior competitor E. B and C are true 32. The competitive exclusion principle states that: A. interference competition will prevail over exploitation B. larger species tend to replace smaller species C. species with the higher productivity will win D. no two species can exist for long on the same limiting resource 33. Which is true regarding succession? A. Climax species permit early-successional species to colonize the disturbed environment B. Pioneer species may modify the environment and permit the establishment of late-successional species. C. Via inhibition, predator species permit several prey species to coexist D. During early stages, competition is more intense than in later stages E. B and D are both true.

29 to 33 please
Answer #1

Answer. 29 D) both A and C are correct because in density dependent factor are those factor where the growth or population size vary on the population density.

Answer 30. D = 1- {\sumn(n-1)/N(N-1)}

Where n = total number of organism of a particular species

N = total number of organism of all species.

N(N-1) = 100 (99) = 9900

\sum n(n-1) = 3770

D = 1- ( 3770/9900) = 1- 0.387 = 0.613.

Answer 31. Character displacement is the result of interspecific competition

Answer 32 D)/ no two species can exist for long on the same limiting resource.

Answer 33. Pioneer species may modify the environment and permit the establishment of later successional species.